5. How many milliliters of a 2.0 M CuSO4 solution are needed to prepare 0.250 L of
0.400 M CuSO_4?
6. What is the molarity of a solution in which 111 g sodium hydrogen carbonate is
dissolved in a 20.0 L solution?
7. A solution of hydrogen peroxide is 13.7% by mass. What is the molarity of the
solution? Assume that the solution has a density of 1.01 g/mL.
8. A solution is prepared by dissolving 19.2 g of glucose in 205 g of water. The final
volume of the solution is 214 ml. Find the concentration of the solution in units of
molality.
1
Expert's answer
2021-09-15T02:40:32-0400

250 ml × 0.4 = 100ml

= 100/2 = 50ml


Molar mass of sodium hydrogen carbonate = 84.007

Molarity = 111/84.007 = 1.32 mol

1.32/20

= 0.066mol/L